QUESTION 1:
Suppose that the current spot exchange rate is GBP1= €1.50 and the one-year forward exchange rate is GBP1=€1.60.

One-year interest rate is 5.4% in euros and 5.2% in pounds.

If you have EUR1,000,000, what is the Covered Interest arbitrage profit in EUR?


QUESTION 2:

Suppose that the current spot exchange rate is GBP1= €1.50 and the one-year forward exchange rate is GBP1=€1.60.
One-year interest rate is 5.4% in euros and 5.2% in pounds.
If you conduct covered interest arbitrage using EUR 25,000,000, which of the following will happen in the market?

A. EUR will depreciate in spot market
B. GBP will appreciate in forward market
C. Interest rate in EUR will decrease
D. Interest rate in GBP will increase
E. None of the above is correct


QUESTION 3:

You have following information:
Spot exchange rate: CAD1=USD0.80
1-year forward rate : CAD1=USD0.79
1-year Canadian interest rate 4%
1-year U.S. interest rate 2.5%

What is the rate of return (in percent) for covered interest arbitrage.

QUESTION 4:

You have following information:
Spot exchange rate: CAD1=USD0.80
1-year forward rate : CAD1=USD0.79
1-year Canadian interest rate 4%
1-year U.S. interest rate 2.5%

When you conduct Covered Interest Arbitrage, which of the following is correct

A. CAD will depreciate in forward market
B. Interest rate in US will drop
C. Interest rate in Canada will increase
D. CAD will depreciate in spot market


QUESTION 5:

Spot rate is: USD1 = HKD7.5000
1-year Forward rate is: USD1 = HKD7.6875

What is the forward rate premium/discount for USD?

QUESTION 6:

Spot rate is: USD1 = HKD7.5000
1-year Forward rate is: USD1 = HKD7.6875
What is the forward rate premium/discount for HKD?
____percent (dont include % sign)
Answer #1

(QUESTION 1)

Here, if covered interest arbitrage relationship holds,

1 + Euro-zone interest rate = (Forward rate / Spot rate) x (1 + British interest rate)

1 + 0.054 = (1.6 / 1.5) x (1 + 0.052)

1.054 = 1.0667 x 1.052

1.054 = 1.1221

Since left-hand side is lower than right-hand side, it means that forward GBP is overvalued. So, investor will borrow Euro.

Therefore, 1,000,000 euros are borrowed at 5.4%.

After 1 year, (Amount + Principal) of this borrowing in Euro = 1,000,000 x 1.054 = 1,054,000

Using this, GBP purchased = 1,000,000 / 1.5 = 666,667

When GBP 666,667 is invested at 5.2%, Amount received after 1 year = 666,667 x 1.052 = GBP 701,333.68

After 1 year, GBP 701,333.68 is sold (using exchange rate of 1.6) for (701,333.68 x 1.6) = Euro 1,122,133.89

Arbitrage profit = (1,122,133.89 - 1,054,000) euro = 68,133.89 euro

NOTE: As per Chegg Answering Policy, 1st question has been answered.